OpenStudy (anonymous):

How would you solve this? c+ (4-3c)-2=0

OpenStudy (saifoo.khan):

c + 4 - 3c - 2= 0 Add like terms, -2c + 2 = 0 -2c = -2 c = -2/-2 c = 1

OpenStudy (anonymous):

where did the -2 come from?

OpenStudy (saifoo.khan):

When +2 is shifted on the other side, it changes it's sign to -ve.